From OpenWetWare
Jump to: navigation, search
Owwnotebook icon.png Our Field of Dreams by "Emblazon" Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

Entry title

Our Field of Dreams

Tiffany ran a formeldehyde agarose gel using protocols from Dr Schwekendiek's lab. The gel showed that there was no RNA in the sample. Due to some errors in the process, the gel will be reran on Thursday.

Rajiv and Melissa started the PCR for the DNA sequence with the primers. The primers were diluted to a 10mM stock and stored in -20C. The Annealing temperature was decided upon from the manufactor's suggested temperator on IDT under Oilgo Analyzer [1]. The suggested annealing temperature for the Forward primer CAT GTC CAT CTT CCT CAT CGC AAC C was 59.9C. And the suggested annealing temperature for the reverse primer CTC CTT TAT CAA ACA ACC CCA TAC G was 55.8C. The PCR was completed using both DNA samples that were stored. We used a pre-made Master mix of dNTP, Taq and 10x Buffer. Then we made a premix of the master mix and primers at 10X which will make a Total of 10 PCR reactions. We used two controls a negative control without any DNA to test if the DNA is the possible problem for no results. And a positive control using a DNA and primer set that is known to work under these conditions, this will show if the Master mix is still viable or not. The DNA was tested using two different concentrations one at 2 micro-liters and 8.4 micro-liters. The PCR was run on a gradient at two temperatures one at 53.7 C and the other at 58.0 C for a total of 30 cycles. What we learned from this experiment is to use one DNA sample when running a PCR. Thrusday we will run the results using electrophoresis, if they worked then we know the annealing temperature to use. If the experiment did not work then we will re-due the experiment with one change using one DNA sample. And if we have time Possibly Redue the RNA extraction due to poor results.

Diwash, Angie, Binu and Sushma made the glycerol stock as well as did the plasmid extraction from the cultures prepared the day before by Angie and Binu. The protocol followed was the Miniprep/GET buffer [2] which was already done once by Angie and Diwash on 25th September. The plasmid was suspended in TE buffer and kept at -20 C, so that the restriction digestion and running on the gel can be done on Thursday. Sushma and Binu also grew new cultures of pSB1A3 and pSB1A7, with which we can work on if the former plasmids fail to work.